OpenStudy (anonymous):

Suppose that the equation, y = 600(1.05)x, represents the average cost of an education at a public college, where x stands for the number of years after 1990. Predict the average cost for 2010 (x = 20), to the nearest dollar. $1,255 $1,592 $2,350 $2,864

OpenStudy (callisto):

Put x =20 into y=600(1.05)^x y = 600 (1.05)^(20) = ... ?

OpenStudy (anonymous):

1,592???

OpenStudy (callisto):

Yes, I think!